# Metrology of microwave fields based on trap-loss spectroscopy with cold Rydberg atoms

### v1

@@ I Introduction
In this Letter, we report on a new method for measuring MW fields with cold Rydberg atoms, based on trap-loss spectroscopy in a magneto-optical trap (MOT). More specifically, we achieve an effective two-photon coupling between the ground state and a Rydberg state, with lasers largely detuned from the intermediate state as in reference [17], in a regime where the extra losses induced by blackbody ionization result in a large fluorescence change as the two-photon laser detuning is scanned across the ground-to-Rydberg atomic transition. This method is simple to implement and in particular it does not require complex detection schemes such as field ionization. Moreover, because the effective two-photon Rabi frequency of the laser coupling (on the order of 9 kHz) is much weaker than the typical MW Rabi frequency to be measured, we realize a quasi-ideal Autler-Townes configuration [26], where the eigenstates resulting from the coupling of two Rydberg states by the MW field are only marginally perturbed by the laser fields. This simple effective two-level configuration is favorable in the context of sensing, and allows in principle the determination of the amplitude and the frequency of the MW field simultaneously from the frequencies of the two Autler-Townes spectral lines, without the need for a local oscillator. We expect this new measurement technique to be particularly well adapted to applications where accuracy, long-term stability and good resolution at large integration times are needed (see the discussion section). The stability of our experimental setup allows to reach a spectral resolution of 20 kHz, corresponding to an equivalent field amplitude of 5 $\mu$V.cm-1, without any observed drift over our 2 hours acquisition period. In the presence of a resonant MW field, we clearly resolve the fluctuations of our MW source.
@@ II Results
Our experimental setup is sketched on figure 1(a). We typically prepare about $10^{7}$ 87Rb atoms in a MOT. Throughout the measurement process, we leave a repumper beam on to avoid the accumulation of atoms in the hyperfine ground state |$\ket{5S_{1/2},F=1}$. We couple the hyperfine state |$\ket{5S_{1/2},F=2}$ to the Rydberg state |$\ket{61S_{1/2}}$ using a two-photon transition with a large detuning $\Delta\sim 2\pi\times 500$ MHz from the intermediate state |$\ket{5P_{3/2},F^{\prime}=3}$ of natural linewidth $2\pi\times 6.065$ MHz (see figure 1(b) and experimental details in subsection IV.1). As we scan the two-photon detuning across the resonance between |$\ket{5S_{1/2},F=2}$ and |$\ket{61S_{1/2}}$, we observe a drop in the fluorescence signal emitted by the atoms, as shown in figure 2(a), allowing to precisely estimate the transition frequency. This technique, known as trap loss spectroscopy, has been used before in various contexts including the study of cold cesium molecules [28, 29] or Rydberg series in cesium [30, 31], strontium [32] and ytterbium [33]. In our case, the atom loss on resonance can be well captured by the following rate equations model [34] taking into account losses from blackbody ionization in the Rydberg state [35]:
@@
So far, we have neglected the effect of the MOT cooling light on the spectrum of figure 2(a). It can be taken into account by a more complete model, based on optical Bloch equations (see section IV.3). This model predicts two transitions of different strengths and frequencies between the ground and the Rydberg states, which can be interpreted as an Autler-Townes doublet resulting from the mixing of |$\ket{5S_{1/2},F=2}$ and |$\ket{5P_{3/2},F^{\prime}=3}$ by the MOT cooling light. As the second transition has a much lower probability, it results in a very small spectral feature (the latter is barely visible on figure 2(a), but its presence is revealed when the numerical model is superposed to the same experimental data, as shown on figure 6 in section IV.3) which we neglect in the rest of this manuscript. More importantly, the MOT cooling light shifts the frequency of the main transition by about $7.6$ MHz, which we take into account in our data analysis. It should be noticed that a similar situation was studied in details in reference [31], where the coupling between ground and Rydberg states was performed by a direct single-photon excitation and where the experimental parameters led to a trap-loss spectrum with more symmetrical Autler-Townes doublet amplitudes.
Figure 2: MOT fluorescence as a function of the two-photon detuning for various applied MW powers ($\omega_{2ph}$ being defined as the sum of the angular frequencies of the two Rydberg lasers). The positions of the two dips on figures 2(b) and 2(c) correspond to $\omega_{2ph}=\omega_{\pm}$ as defined by equation (2). Black dots are experimental data, and red curves were derived from the rate equations model described in section IV.2. The scanning rate is 2 MHz/s.
Another significant effect arises from the finite response time of the MOT compared to the scanning time (more precisely the width of the spectral profile divided by the scanning rate). This leads to asymmetric spectral profiles depending on the speed of the frequency scans. To mitigate this effect, we adopt a moderate scanning rate of 2 MHz/s (unless otherwise specified) and we use asymmetric gaussian functions to fit the spectral profiles. For each measurement, we estimate the transition frequency by averaging between the two opposite scanning directions (see section IV.4 for more details).
Throughout this manuscript, we define the origin of the frequency axis as our best estimate (see II.4) of the position of the fluorescence dip while probing the |$\ket{5S_{1/2},F=1}\leftrightarrow\ket{61S_{1/2}}$ transition in the absence of applied MW field, denoted $\omega_{0}$. In particular, $\omega_{0}$ includes light shifts on the ground state induced by the MOT cooling light, as discussed earlier in this section.
@@ II.2 Autler-Townes doublet in the presence of MW
We now apply the MW signal to be measured, with a frequency at 15.973 GHz, close to resonance with the |$\ket{61S_{1/2}}$ to |$\ket{61P_{1/2}}$ transition [36]. Typically (see figure 3 and section II.3), the measured MW Rabi frequency is larger than 10MHz, which is much larger than the two-photon Rabi frequency between |$\ket{5S_{1/2},F=2}$ and |$\ket{61S_{1/2}}$ of about $9$ kHz. Consequently, this system is very close to the textbook Autler-Townes doublet configuration, where a coupled two-level system is probed by a weak field, resulting in two spectral lines at frequencies:
@@
where $\Omega_{MW}$ and $\Delta_{MW}$ are respectively the Rabi frequency and the detuning of the MW field acting on the |$\ket{61S_{1/2}}\leftrightarrow\ket{61P_{1/2}}$ transition. Examples of such Autler-Townes doublets are shown on fig.2(b) and 2(c). On resonance ($\Delta_{MW}=0$), the frequency difference between the two spectral profiles $\delta\omega=\omega_{+}-\omega_{-}$ provides a measurement of the MW Rabi frequency $\Omega_{MW}$, which is proportional to the amplitude of the MW electric field. Interestingly, the center frequency $\bar{\omega}=(\omega_{+}+\omega_{-})/2$ can be used to measure the frequency of the MW field, as will be discussed in section II.5.
@@
II.3 Linearity and measurement range
Figure 3: Measured Autler-Townes splitting as a function of the square root of the MW power at 15.973 GHz sent to the antenna. The red curve is a linear fit to the data, yielding an intercept value of $1.6\pm 0.5$ MHz.
The measured Autler-Townes splitting $\delta\omega$ as a function of the applied MW field is shown on figure 3. It was obtained by sending on the atoms a MW field at $15.973$ GHz, resonant with the |$\ket{61S_{1/2}}\leftrightarrow\ket{61P_{1/2}}$ transition ($|\Delta_{MW}|/\Omega_{MW}\ll 1$). For each value of the power sent at the MW horn input, each data point was obtained by averaging over 10 successive pair of scans of opposite directions. The error bars for the $y$ axis correspond to the standard deviation of the 10 measurements. The error bars in $x$ correspond to an estimated 0.1 dB fluctuation on the MW power delivered to the atoms.
As can be seen on the lower part of figure 3, the relative deviation between the experimental points and the fit is on the order of $2\,\%$ over the whole measurement range. We attribute this small discrepancy mostly to the uncertainty on the applied MW power and to the residual frequency, amplitude and polarization noise of the MOT and Rydberg lasers (see II.4 and IV.1).
The intercept value resulting from the linear fit is $1.6\pm 0.5$ MHz, corresponding to a bias in the measurement which we attribute to our fitting protocol: when the two Autler-Townes traces start to overlap, the dynamics of the MOT results in an overall trace which slightly differs from the sum of two asymetric gaussians that we use for fitting the data. This bias could be mitigated by fitting with a trace resulting from a time-dependent simulation of the rate equations as described in the supplementary sections IV.2 and IV.4.
@@ II.3 Linearity and measurement range
II.4 Resolution, sensitivity and long-term stability
In order to evaluate the metrological performance of this measurement technique, we first record, without any applied MW signal, 57 pairs of trap-loss spectra of the ground to |$\ket{61S_{1/2}}$ transition within identical experimental conditions. For each scan, the two photon laser frequency is swept at 2 MHz/s from $-50$ MHz to $+50$ MHz across the reference frequency $\omega_{0}/2\pi$ and back, thus setting the total measurement time to 5700 s. The overlapping Allan deviation of the extracted transition frequency is depicted in figure 4 (blue points).
At short-term this Allan deviation aligns well with a trend curve of $1$MHz$/\sqrt{\tau}$ (black line), where $\tau$ is the integration time. This $1/\sqrt{\tau}$ scaling indicates that the short-term sensitivity is limited by white noise. In particular for $\tau=100$ s the shot-to-shot fluctuation of the transition frequency is well-captured by the relative amplitude noise of 5 to $10\,\%$ on the fluorescence signals (see figure 2(a)), which mostly results from the few MHz linewidth of the cooling laser light. For higher values of $\tau$, the Allan deviation demonstrates the capability of our setup to resolve frequency changes as low as 20 kHz for an integration time of about 2500 s.
These numbers, measured in the absence of applied MW signal, can be used to extrapolate the measurement performance with the assumption that the MW induced Autler-Townes splitting can be extracted at the same level than the bare two-photon transition frequency. This leads to an inferred short-term sensitivity of 247 $\mu$V.cm-1.Hz-1/2 and a resolution of 5 $\mu$V/cm based on the theoretical dipole matrix element of the |$\ket{61S_{1/2}}\leftrightarrow\ket{61P_{1/2}}$ transition. It should be pointed out that the measurement method described here is not optimal for maximizing the sensitivity. Indeed, a more efficient approach would be to operate the sensor only at the points where the slope of the fluorescence signal is maximal.
@@ II.4 Resolution, sensitivity and long-term stability
Figure 4: Overlapping Allan deviation of the transition frequency in the absence of MW (blue) and of the Autler-Townes splitting frequency when a MW signal at 15.973 GHz and about -23 dBm is sent to the antenna (green). The black line is a guide to the eye showing a $1/\sqrt{\tau}$ trend of the data in blue.
II.5 Experimental results in the case of a non-resonant MW field
Figure 5: (a) Measured Autler-Townes splitting versus MW detuning (relatively to the theoretical transition frequency of 15.973 GHz) for 3 different powers sent to the antenna (P = -20 dBm, red ; P = -17 dBm, green ; P = -14 dBm, blue). (b) Measured Autler-Townes doublet center frequency versus MW detuning from the same dataset. In both plots, the data (color circles) are adjusted by the model of equations (3) and (4) with parameters $\Delta_{0}=-2\pi\times 1.36\,$MHz and $\Omega_{MW}=2\pi\times 47\,$MHz (red dashed lines); $2\pi\times 66.5\,$MHz (green dashed lines); $2\pi\times 94.1\,$MHz (blue dashed line). The fact that the three lines on figure (b) have different offsets is a consequence of light shifts, see appendix IV.5 for more details.
@@ II.5 Experimental results in the case of a non-resonant MW field
where $\delta_{1,2}^{LS}$ are the overall light shifts on the |$\ket{61S_{1/2}}$ and |$\ket{61P_{1/2}}$ states respectively. The theoretical calculation and experimental measurement of $\delta_{1,2}^{LS}$ is described in appendix IV.5. To give a sense of scale, for a MW power corresponding to $\Omega_{MW}=2\pi\times 100\;\textrm{MHz}$, we find $\delta_{1}^{LS}\simeq 2\pi\times(-7.4)\;\textrm{MHz}$ and $\delta_{2}^{LS}\simeq 2\pi\times(-0.4)\;\textrm{MHz}$ respectively.
@@
where $\omega_{MW}$ is the applied MW frequency, $\omega_{ARC}=2\pi\times 15.973$ GHz is the atomic transition frequency from ARC Rydberg calculator [36] and $\Delta_{0}\simeq-2\pi\times 1.36$ MHz according to the adjustment of the data by the model as shown on figure 5. We attribute the origin of the residual discrepancy between the data and the model (up to a few MHz at most) to be mainly due to fluctuations in the applied MW power. This is supported by the fact that the $\delta\omega$ curves (figure 5(a)), which have a much stronger dependence in the MW power than $\bar{\omega}$ according to the theory, also have a much higher level of discrepancy. Other possible causes for the observed discrepancy include: the deterioration of the signal to noise ratio for one of the two spectral profiles as the value of the detuning increases and becomes comparable with $\Omega_{MW}$; the bias induced by our fitting protocol as discussed in section II.3; the effects of the finite response time of the MOT (section IV.4), which might be enhanced when the amplitudes of the two spectral profiles become significantly different; the uncertainties in the estimation of the dipole matrix elements for the light shift calculation, which might become more and more important for increasing values of the MW power.
@@
III Discussion
In this paper, we have demonstrated and characterized a new method to measure MW fields with Rydberg atoms based on trap-loss spectroscopy in a magneto-optical trap. This method is particularly simple as the detection scheme relies on fluorescence measurements only. By using a two-photon transition highly-detuned from the intermediate state, we realize a situation where the frequencies of the spectral lines are well-described by a coupled two-level system, which is particularly favorable for the linearity of the sensor in the resonant case. In the non-resonant case, this simple two-level system behavior allows in principle to reconstruct both the amplitude and the frequency of the applied MW field from the Autler-Townes splitting frequency and doublet center frequency, provided the light shifts induced by the MW field are taken into account. The maximum measurement range that we achieved is already much higher than for electromagnetically-induced absorption in cold atoms [17], and could be possibly much larger if the coupling (blue) laser was scanned instead of the probe (infrared) laser, which could be technically facilitated by the use of the inverted scheme [37] where the coupling laser is in the infrared domain (1010nm). Future directions to improve the metrological performance of our experimental setup include a better control of the intensity of the cooling beams, Rydberg beams and their associated polarizations, as well as a better control of the MW power effectively sent onto the atoms. In the longer term, one could get rid of the MOT broadening by adapting the setup to perform pulsed MOT operation, optical molasses or even dipole traps.

### v2

@@ I Introduction
In this Letter, we report on a new method for measuring MW fields with cold Rydberg atoms, based on trap-loss spectroscopy in a magneto-optical trap (MOT). More specifically, we achieve an effective two-photon coupling between the ground state and a Rydberg state, with lasers largely detuned from the intermediate state as in reference [17], in a regime where the extra losses induced by blackbody ionization result in a large fluorescence change as the two-photon laser detuning is scanned across the ground-to-Rydberg atomic transition. This method is simple to implement and in particular it does not require complex detection schemes such as field ionization. Moreover, because the effective two-photon Rabi frequency of the laser coupling (on the order of 9 kHz) is much weaker than the typical MW Rabi frequency to be measured, we realize a quasi-ideal Autler-Townes configuration [26], where the eigenstates resulting from the coupling of two Rydberg states by the MW field are only marginally perturbed by the laser fields. This simple effective two-level configuration is favorable in the context of sensing, and allows in principle the determination of the amplitude and the frequency of the MW field simultaneously from the frequencies of the two Autler-Townes spectral lines, without the need for a local oscillator. We expect this new measurement technique to be particularly well adapted to applications where accuracy, long-term stability and good resolution at large integration times are needed. The stability of our experimental setup allows to reach a spectral resolution of 20 kHz, corresponding to an equivalent field amplitude of 5 $\mu$V.cm-1, without any observed drift over our 2 hours acquisition period. In the presence of a resonant MW field, we clearly resolve the fluctuations of our MW source.
@@ II Measurement principle
Our experimental setup is sketched on figure 1(a). We typically prepare about $10^{7}$ 87Rb atoms in a MOT. Throughout the measurement process, we leave a repumper beam on to avoid the accumulation of atoms in the hyperfine ground state |$\ket{5S_{1/2},F=1}$. We couple the hyperfine state |$\ket{5S_{1/2},F=2}$ to the Rydberg state |$\ket{61S_{1/2}}$ using a two-photon transition with a large detuning $\Delta\sim 2\pi\times 500$ MHz from the intermediate state |$\ket{5P_{3/2},F^{\prime}=3}$ of natural linewidth $2\pi\times 6.065$ MHz (see figure 1(b) and experimental details in subsection VI.1). As we scan the two-photon detuning across the resonance between |$\ket{5S_{1/2},F=2}$ and |$\ket{61S_{1/2}}$, we observe a drop in the fluorescence signal emitted by the atoms, as shown in figure 2(a), allowing to precisely estimate the transition frequency. This technique, known as trap loss spectroscopy, has been used before in various contexts including the study of cold cesium molecules [27, 28] or Rydberg series in cesium [29, 30], strontium [31] and ytterbium [32]. In our case, the atom loss on resonance can be well captured by the following rate equations model [33] taking into account losses from blackbody ionization in the Rydberg state [34]:
@@
So far, we have neglected the effect of the MOT cooling light on the spectrum of figure 2(a). It can be taken into account by a more complete model, based on optical Bloch equations (see section VI.3). This model predicts two transitions of different strengths and frequencies between the ground and the Rydberg states, which can be interpreted as an Autler-Townes doublet resulting from the mixing of |$\ket{5S_{1/2},F=2}$ and |$\ket{5P_{3/2},F^{\prime}=3}$ by the MOT cooling light. As the second transition has a much lower probability, it results in a very small spectral feature (the latter is barely visible on figure 2(a), but its presence is revealed when the numerical model is superposed to the same experimental data, as shown on figure 6 in section VI.3) which we neglect in the rest of this manuscript. More importantly, the MOT cooling light shifts the frequency of the main transition by about $7.6$ MHz, which we take into account in our data analysis. It should be noticed that a similar situation was studied in details in reference [30], where the coupling between ground and Rydberg states was performed by a direct single-photon excitation and where the experimental parameters led to a trap-loss spectrum with more symmetrical Autler-Townes doublet amplitudes.
Figure 2: MOT fluorescence as a function of the two-photon detuning for various applied MW powers ($\omega_{2ph}$ being defined as the sum of the angular frequencies of the two Rydberg lasers). The positions of the two dips on figures 2(b) and 2(c) correspond to $\omega_{2ph}=\omega_{\pm}$ as defined by equation (2). Black dots are experimental data, and red curves were derived from the rate equations model described in section VI.2. The scanning rate is 2 MHz/s.
Another significant effect arises from the finite response time of the MOT compared to the scanning time (more precisely the width of the spectral profile divided by the scanning rate). This leads to asymmetric spectral profiles depending on the speed of the frequency scans. To mitigate this effect, we adopt a moderate scanning rate of 2 MHz/s (unless otherwise specified) and we use asymmetric gaussian functions to fit the spectral profiles. For each measurement, we estimate the transition frequency by averaging between the two opposite scanning directions (see section VI.4 for more details).
Throughout this manuscript, we define the origin of the frequency axis as our best estimate (see III.2) of the position of the fluorescence dip while probing the |$\ket{5S_{1/2},F=1}\leftrightarrow\ket{61S_{1/2}}$ transition in the absence of applied MW field, denoted $\omega_{0}$. In particular, $\omega_{0}$ includes light shifts on the ground state induced by the MOT cooling light, as discussed earlier in this section.
@@ II.2 Autler-Townes doublet in the presence of MW
We now apply the MW signal to be measured, with a frequency at 15.973 GHz, close to resonance with the |$\ket{61S_{1/2}}$ to |$\ket{61P_{1/2}}$ transition [35]. Typically (see figure 3 and section III.1), the measured MW Rabi frequency is larger than 10MHz, which is much larger than the two-photon Rabi frequency between |$\ket{5S_{1/2},F=2}$ and |$\ket{61S_{1/2}}$ of about $9$ kHz. Consequently, this system is very close to the textbook Autler-Townes doublet configuration, where a coupled two-level system is probed by a weak field, resulting in two spectral lines at frequencies:
@@
where $\Omega_{MW}$ and $\Delta_{MW}$ are respectively the Rabi frequency and the detuning of the MW field acting on the |$\ket{61S_{1/2}}\leftrightarrow\ket{61P_{1/2}}$ transition. Examples of such Autler-Townes doublets are shown on fig.2(b) and 2(c). On resonance ($\Delta_{MW}=0$), the frequency difference between the two spectral profiles $\delta\omega=\omega_{+}-\omega_{-}$ provides a measurement of the MW Rabi frequency $\Omega_{MW}$, which is proportional to the amplitude of the MW electric field. Interestingly, the center frequency $\bar{\omega}=(\omega_{+}+\omega_{-})/2$ can be used to measure the frequency of the MW field, as will be discussed in section IV.
@@
III Performance evaluation in the case of a resonant MW field
III.1 Linearity and measurement range
Figure 3: Measured Autler-Townes splitting as a function of the square root of the MW power at 15.973 GHz sent to the antenna. The error bars on the x-axis correspond to an estimated uncertainty of 0.1 dB on the applied MW power. The error bars on the y-axis reflect the 2-$\sigma$ standard deviation of 10 successive measurements for each point. Their amplitude is typically smaller than 0.5 MHz. The red curve is a linear fit to the data, yielding an intercept value of $1.6\pm 0.5$ MHz.
The measured Autler-Townes splitting $\delta\omega$ as a function of the applied MW field is shown on figure 3. It was obtained by sending on the atoms a MW field at $15.973$ GHz, resonant with the |$\ket{61S_{1/2}}\leftrightarrow\ket{61P_{1/2}}$ transition ($|\Delta_{MW}|/\Omega_{MW}\ll 1$). For each value of the power sent at the MW horn input, each data point was obtained by averaging over 10 successive pair of scans of opposite directions. The error bars for the $y$ axis correspond to the 2-$\sigma$ standard deviation of the 10 measurements. The error bars in $x$ correspond to an estimated 0.1 dB fluctuation on the MW power delivered to the atoms.
As can be seen on the lower part of figure 3, the relative deviation between the experimental points and the fit is on the order of $2\,\%$ over the whole measurement range. We attribute this small discrepancy mostly to the uncertainty on the applied MW power and to the residual frequency, amplitude and polarization noise of the MOT and Rydberg lasers (see III.2 and VI.1).
The intercept value resulting from the linear fit is $1.6\pm 0.5$ MHz, corresponding to a bias in the measurement which we attribute to our fitting protocol: when the two Autler-Townes traces start to overlap, the dynamics of the MOT results in an overall trace which slightly differs from the sum of two asymetric gaussians that we use for fitting the data. This bias could be mitigated by fitting with a trace resulting from a time-dependent simulation of the rate equations as described in the supplementary sections VI.2 and VI.4.
@@ III.1 Linearity and measurement range
III.2 Resolution, sensitivity and long-term stability
In order to evaluate the metrological performance of this measurement technique, we first record, without any applied MW signal, 57 pairs of trap-loss spectra of the ground to |$\ket{61S_{1/2}}$ transition within identical experimental conditions. For each scan, the two photon laser frequency is swept at 2 MHz/s from $-50$ MHz to $+50$ MHz across the reference frequency $\omega_{0}/2\pi$ and back, thus setting the total measurement time to 5700 s.
The overlapping Allan deviation of the extracted transition frequency $\sigma_{\omega}$ as a function of the integration time $\tau$ is depicted in figure 4 (blue points). We use this Allan deviation curve, taken in the absence of MW, to infer the intrinsic stability of our experiment, and extrapolate the measurement performance with the assumption that the MW induced Autler-Townes splitting can be extracted at the same level than the bare two-photon transition frequency.
[…]
For increasing values of $\tau$, the Allan deviation curve demonstrates the capability of our setup to resolve increasingly smaller frequency changes. We define the resolution as the smallest value of the Allan deviation measured with our setup, on the order of 20 kHz for an integration time of about 2500 s as can be seen on figure 4. Once again, this frequency resolution can be converted into an electric field resolution using the appropriate transition dipole matrix element, yielding $\simeq$ 5 $\mu$V/cm.
We expect the long term stability of our setup to be ultimately limited by variations of the cooling light intensity, which creates lightshifts on the order of several MHz on the |$\ket{5S_{1/2}}$ ground state as discussed in sections II.1 and VI.3 (figure 6). Variations in the polarization of the cooling light and in the intensity and polarization of the Rydberg lasers might also affect the long term stability.
@@ III.2 Resolution, sensitivity and long-term stability
Figure 4: Overlapping Allan deviation of the transition frequency $\sigma_{\omega}$ in the absence of MW (blue) and of the Autler-Townes splitting frequency $\sigma_{\delta\omega}$ when a MW signal at 15.973 GHz and -23 dBm is sent to the antenna (green). The black line is a guide to the eye showing a $1/\sqrt{\tau}$ trend of the data in blue.
IV Experimental results in the case of a non-resonant MW field
Figure 5: (a) Measured Autler-Townes splitting versus MW detuning (relatively to the theoretical transition frequency of 15.973 GHz) for 3 different powers sent to the antenna (PMW = -20 dBm, red ; PMW = -17 dBm, green ; PMW = -14 dBm, blue). (b) Measured Autler-Townes doublet center frequency versus MW detuning from the same dataset. In both plots, the data (color circles) are adjusted by the model of equations (3) and (4) with the following procedure: for an initial value of $\Delta_{0}$ we perform three independent adjustments for the datasets associated to -20 dBm, -17 dBm and -14 dBm respectively to find three values of $\Omega_{MW}$. We then iterate on the value of $\Delta_{0}$ and repeat the procedure until the distance between the fits and the datasets is minimized. This procedure yields to the following values: $\Delta_{0}=-2\pi\times 1.36\,$MHz and $\Omega_{MW}=2\pi\times 47\,$MHz (red dashed lines, PMW = -20 dBm); $2\pi\times 66.5\,$MHz (green dashed lines, PMW = -17 dBm); $2\pi\times 94.1\,$MHz (blue dashed line, PMW = -14 dBm). The fact that the three lines on figure (b) have different offsets is a consequence of light shifts, see appendix VI.5 for more details.
@@ IV Experimental results in the case of a non-resonant MW field
where $\delta_{1,2}^{LS}$ are the overall light shifts on the |$\ket{61S_{1/2}}$ and |$\ket{61P_{1/2}}$ states respectively. The theoretical calculation and experimental measurement of $\delta_{1,2}^{LS}$ is described in appendix VI.5. To give a sense of scale, for a MW power corresponding to $\Omega_{MW}=2\pi\times 100\;\textrm{MHz}$, we find $\delta_{1}^{LS}\simeq 2\pi\times(-7.4)\;\textrm{MHz}$ and $\delta_{2}^{LS}\simeq 2\pi\times(-0.4)\;\textrm{MHz}$ respectively.
@@
where $\omega_{MW}$ is the applied MW frequency, $\omega_{ARC}=2\pi\times 15.973$ GHz is the atomic transition frequency from ARC Rydberg calculator [35] and $\Delta_{0}\simeq-2\pi\times 1.36$ MHz according to the adjustment of the data by the model as shown on figure 5. We attribute the origin of the residual discrepancy between the data and the model (up to a few MHz at most) to be mainly due to fluctuations in the applied MW power. This is supported by the fact that the $\delta\omega$ curves (figure 5), which have a much stronger dependence in the MW power than $\bar{\omega}$ according to the theory, also have a much higher level of discrepancy. Other possible causes for the observed discrepancy include: the deterioration of the signal to noise ratio for one of the two spectral profiles as the value of the detuning increases and becomes comparable with $\Omega_{MW}$; the bias induced by our fitting protocol as discussed in section III.1; the effects of the finite response time of the MOT (section VI.4), which might be enhanced when the amplitudes of the two spectral profiles become significantly different; the uncertainties in the estimation of the dipole matrix elements for the light shift calculation, which might become more and more important for increasing values of the MW power.
@@
V Conclusion
In this paper, we have demonstrated and characterized a new method to measure MW fields with Rydberg atoms based on trap-loss spectroscopy in a magneto-optical trap. This method is particularly simple as the detection scheme relies on fluorescence measurements only. By using a two-photon transition highly-detuned from the intermediate state, we realize a situation where the frequencies of the spectral lines are well-described by a coupled two-level system, which is particularly favorable for the linearity of the sensor in the resonant case. In the non-resonant case, this simple two-level system behavior allows in principle to reconstruct both the amplitude and the frequency of the applied MW field from the Autler-Townes splitting frequency and doublet center frequency, provided the light shifts induced by the MW field are taken into account. The maximum range that we achieved for the measurement of the microwave power, on the order of 160 MHz, is already much higher than for electromagnetically-induced absorption in cold atoms [17], and could be possibly much larger if the coupling (blue) laser was scanned instead of the probe (infrared) laser, which could be technically facilitated by the use of the inverted scheme [36] where the coupling laser is in the infrared domain (1010nm). The maximum range that we achieved for the measurement of the microwave frequency is on the order of ±$\pm 50$ MHz around the resonance frequency, typically set by the value of the MW Rabi frequency on resonance. Future directions to improve the metrological performance of our experimental setup include a better control of the intensity of the cooling beams, Rydberg beams and their associated polarizations, as well as a better control of the MW power effectively sent onto the atoms. In the longer term, one could get rid of the MOT broadening by adapting the setup to perform pulsed MOT operation, optical molasses or even dipole traps.
In comparison with state-of-the-art techniques based on Rydberg EIT in room-temperature vapor cells, this new approach has several advantages, including reduced Doppler and transit time effects, better long-term stability as the atoms are more isolated from their environment, and the possibility to measure simultaneously the amplitude and the frequency of the electric field. On the other hand, the time needed to perform a single measurement with this new approach is much higher than for vapor cells as it is limited by the response time of the magneto-optical trap, on the order of one second in the present work.